# 1.3in1.3in1in1in

### v1

@@ Abstract.
For a cubic fourfold $Y$ with associated Fano variety of lines $F$, we establish several properties of the finite-degree 16 self-rational map $\psi\colon F\dashrightarrow F$ introduced by Voisin. We begin by analyzing the singularities of the nodal quintic with 16 nodes associated with a general line under the specialization to a line in the branch locus of $\psi$. This approach reveals that the ramification of the natural resolution of indeterminacy of $\psi$ is simple. The main part of the paper focuses on the intriguing interplay between $\psi$ and the fixed locus of the antisymplectic involution on the LLSvS variety $Z$, examined via the degree 6 Voisin map $F\times F\dashrightarrow Z$. As an application, we show that the monodromy of $\psi$ is maximal.
@@ 1. Introduction
Let $Y$ be a smooth cubic fourfold and $F$ be its (Fano) variety of lines. The rich geometry of these varieties has attracted great attention from the mathematical community for several reasons. Notably, the Fano variety $F$ is one of the earliest examples of a locally complete family of projective hyperkähler manifolds, and the binome $Y-F$ between a Fano variety and a hyperkähler manifold has inspired numerous other constructions.
One distinguishing feature of $F$ is the degree 16 self-rational Voisin map $\psi\colon F\dashrightarrow F$. Remarkably, $\psi$ is the only known finite, non-birational self-map defined on a locally complete family of projective hyperkähler manifolds. Finite maps such as $\psi$ are particularly subtle to study as, unlike birational maps, they cannot be detected through their action on the second cohomology group and lack a comprehensive classification result as Hodge Torelli theorem that would allow precise control over their behavior. Since its construction [Voi04], this map has continued to attract significant interest in the mathematical community, as demonstrated by the numerous works on the subject[Ame09, GK23b, GK23a, GK22].
In this paper, we focus on the monodromy of the Voisin map $\psi$. The monodromy group is a discrete invariant that encodes the symmetries and intrinsic geometry of finite maps. Determining the monodromy group of branched coverings is a classical problem, originating with Jordan in the 1870s. This topic was revitalized by Harris, who provided a modern framework by proving that the monodromy and Galois groups coincide, and developed tools to establish when the monodromy group is maximal. Significant progress was made by Vakil, who introduced innovative techniques to study monodromy in Schubert problems on Grassmannians [Vak06].
Building on these developments, we take a step further by investigating linear spaces on a cubic hypersurface of dimension 4. Our main result shows that the monodromy group of the Voisin map $\psi$ is maximal, meaning that it is the full symmetric group on 16 elements. To achieve this, we leverage another Voisin map involving twisted cubics on the cubic fourfold, which unveils a fascinating connection with the fixed locus of the natural anti-symplectic involution on the LLSvS variety.
@@
$\Lambda_{\ell}:=\bigcap_{p\in L}T_{p}Y$
is 2-dimensional and in this case the line is said to be of type I. For special lines, called of type II, the dimension of $\Lambda_{\ell}$ is 3. For the general line we have $\Lambda_{\ell}\cap Y=2\ell+r$ for a line $r\in F$, and one sets $\psi(\ell):=r$. This map has been studied in detail in [Ame09], its indeterminacy locus consists of the lines of type $\operatorname{II}$, which for a general cubic fourfold form a smooth surface $S_{\operatorname{II}}$. Blowing up $S_{\operatorname{II}}$ provides a resolution of the indeterminacy
@@
Gounelas and Kouvidakis recently computed that the restricion of ${\widehat{\psi}}$ to the exceptional divisor $E$ of the blow-up morphism, which coincides with the ramification divisor of ${\widehat{\psi}}$, is birational onto the image and they posed the question of whether the ramification of ${\widehat{\psi}}$ is simple [GK23a, Theorem B, Remark 3.13].
In Section 2 we study the ramification of ${\widehat{\psi}}$ by consider the projection of $Y$ from a general line $R$ in $Y$. Its resolution is a conic bundle over $\mathbb{P}^{3}$ with discriminant locus a quintic surface $S_{r}$ with exactly 16 nodes. These nodes correspond to the preimage of $r$ under ${\widehat{\psi}}$. Nodal quintic surfaces have been classically studied by Beauville and Catanese [Bea80, Cat81] and more recently studied in [Huy24, BFK+24, CicwYCC+24]. Our first attention is to study the singularity of $S_{r}$ for a special line $r$, which lies in the branch locus of ${\widehat{\psi}}$.
@@
Let $r\in F$ be a general line in the branch divisor of ${\widehat{\psi}}$. Then the quintic surface $S_{r}$ has one singularity of type $A_{3}$ corresponding to the unique line $L$ of type $\operatorname{II}$ over $r$.
As a result of our analysis we affirmatively answer Gounelas and Kouvidakis’ question:
Corollary 1.2 (see Corollary 2.7).
The ramification at the general point of the ramification locus of ${\widehat{\psi}}$ is simple.
In Section 3 we shift our focus to the geometry of twisted cubics. With any smooth cubic fourfold $Y$ not containing a plane Lehn, Lehn, Sorger, and van Straten associated an eight-dimensional hyperkähler variety $Z$ parametrizing families of twisted cubics and their flat degenerations. The variety $Z$ is equipped with a natural antisymplectic involution $\tau$ [Leh15]. Its fixed locus is a smooth Lagrangian submanifold with 2 connected components: one is isomorphic to the cubic fourfold $Y$, the other one, $W$, is of general type [FMOS23] and remains somewhat mysterious. Using a the degree 6 rational map $\varphi\colon F\times F\dashrightarrow Z$, constructed by Voisin [Voi16], we offer an alternative description of $W$. We define the variety $P$ as the closure in $F\times F$ of
$\{(\ell_{1},\ell_{2})\in F\times F:\ell_{i}\text{ are of type I,\quad$\ell_{1}\not=\ell_{2}$,\quad and }\quad\psi(\ell_{1})=\psi(\ell_{2})\},$
which is birational to an irreducible component of the self-product of ${\widehat{F}}$ over $F$.
@@
In Section 4 after revising the basic notions of monodromy we tackle the study of ${\widehat{\psi}}$. Despite extensive study and numerous works concerning ${\widehat{\psi}}$, for example about its entropy [Ame09], many of its properties remain elusive. Through an investigation of the restriction of the map $\varphi$ to the variety $P$ we prove the following
@@ Acknowledgments
This project began long time ago, and over the years we benefited from conversations with many people. It is our pleasure to thank everybody who expressed interest and shared their point of view, especially Enrico Fatighenti, Frank Gounelas, Christian Lehn, Emanuele Macrì, Giovanni Mongardi, Alan Thompson and Yilong Zhang.
Franco Giovenzana was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Projektnummer 509501007, and partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC-2020- SyG-854361- HyperK). All authors are members of INdAM GNSAGA.
2. The ramification of the Voisin map is simple
In this section we recall basic facts abour lines on cubic fourfolds and various properties of the Voisin map on $F$. Then we move on to prove that the ramification of the Voisin map is simple.
@@ 2. The ramification of the Voisin map is simple
Given a line $L$, we say that $L$ is a line of type I if the equivalent conditions $(1)-(4)$ hold, whereas if $(5)-(8)$ hold we say that $L$ is a line of type II. We set $\Lambda_{L}:=\cap_{P\in L}T_{P}Y$.
@@
In case $\ell\in F$ is of type $\operatorname{II}$, then the line $\mathscr{G}(L)\subset{\mathbb{P}}(V^{\vee})$ is the projective dual of $\Lambda_{L}\subset{\mathbb{P}}(V)$. In case of a line of type I, $\mathscr{G}(L)$ spans a ${\mathbb{P}}^{2}$ in ${\mathbb{P}}(V^{\vee})$ which is dual to $\Lambda_{L}$.
@@
[Huy23, Remark 2.2.2] Let $Y=V(F)\subset{\mathbb{P}}^{5}$ be a smooth cubic fourfold, let $L$ be a line in $Y$, then $L$ is of type II if and only if the partial derivatives $\partial_{0}F|_{L},...,\partial_{6}F|_{L}$ span a vector space of dimension 2 in $H^{0}(L,\mathcal{O}_{\mathbb{P}^{5}}(2))$.
Voisin observed that if $L$ is a line of type I, then $\Lambda_{L}\cap Y$ is a cubic curve, which, as it contains $L$ with multiplicity 2, consists of $L$ and a residual line $R$. As the general line is of type I, one defines the rational map
$\psi\colon F\dashrightarrow F,\quad\ell\mapsto r.$
This map has been studied in [Voi04] and [Ame09], where it is proven to be finite of degree 16. Lines of type II form a surface $S_{\operatorname{II}}$ in $F$, and blowing up $F$ in this surface resolves the indeterminacy of $\psi$:
@@
The blow up ${\widehat{F}}$ can be identified as the closure of the graph [GK22, Lemma 4.1],[Huy23, Remark 2.2.19].
$F\dashrightarrow\operatorname{Gr}(3,6),\ \ell\mapsto\Lambda_{\ell}.$
Elements in the exceptional locus are then just pairs $(\ell,\Xi)$, where $L$ is a line of type II and $\Xi$ is a projective plane such that $L\subset\Xi\subset\Lambda_{\ell}$. For such 2-dimensional spaces $\Xi$ we have that $\Xi\cap Y=2\ell+r$ and we shall say that $r$ is residual to $\ell$.
@@
[…]
where $\pi_{R}$ is the projection from $R$ and ${\widetilde{Y}}$ is the blow-up of $Y$ in $R$. The morphism ${\widetilde{\pi}}$ is a conic bundle with discriminant a quintic surface $S_{r}$, whose singular locus consists of 16 nodes. The 16 nodes correspond to the preimage of $R$ under ${\widehat{\psi}}$. Indeed, if $p_{i}$ denote the nodes of $S_{R}$ for $i=1,..,16$, then $L_{i}:=p({\widetilde{\pi}}^{-1}(p_{i}))$ are the 16 lines for which $\psi(L_{i})=R$ (see [Huy23, § 6.4.5] for an account on the various results in the literature about this surface).
Let $E$ be the exceptional divisor of the blow-up morphism ${\widehat{F}}\to F$. As $F$ has trivial canonical bundle, the divisor $E$ coincide with the ramification locus of the map ${\widehat{\psi}}$.
Theorem 2.5 (Gounelas-Kouvidakis,[GK22]).
The restriction ${\widehat{\psi}}|_{E}\colon E\to F$ is generically 1-to-1 onto the image.
In other words, over the general point $r$ in the branch divisor of ${\widehat{\psi}}$, there is exactly one point of ramification, i.e. there exists exactly one line $L$ of type II with residual $R$, meaning that ${\widehat{\psi}}(\ell,\langle L,R\rangle)=r$.
@@
Let $r\in F$ be a general line in the branch divisor of ${\widehat{\psi}}$. Then the quintic surface $S_{r}$ has exactly 14 singularities of type $A_{1}$ and exactly 1 singularity of type $A_{3}$ corresponding to the unique line $L$ of type $\operatorname{II}$ with residual line $R$.
@@
Let $r$ be a general line in the branch locus of ${\widehat{\psi}}$ and let $\ell$ be the unique line of type $\operatorname{II}$ with residual $r$. We may assume that $R$ is of type I and that is given as the zero set $V(x_{2},x_{3},x_{4},x_{5})$. Following Clemens and Griffiths [CG72, equation 6.9] we write the equation of the cubic fourfold as
(1) $x_{4}x_{0}^{2}+x_{5}x_{0}x_{1}+x_{3}x_{1}^{2}+x_{0}Q_{0}+x_{1}Q_{1}+P=0,$
@@
We assume that the line $L$ is given by $V(x_{1},x_{2},x_{4},x_{5})$ [^1] Notice that with this assumption the line $L$ is not contained in the plane tangent to $R$ is $\Lambda_{r}=V(x_{3},x_{4},x_{5})$, would have been in contradiction with $R$ being residual to $L$., then $p:=(1:0:0:0:0:0)$ is the intersection of $L$ and $R$, and $q:=(0:0:0:1:0:0)$ is the point of intersection with $\mathbb{P}^{3}$.
@@
The cubic fourfold $Y$ contains the line $L$ if $F(x_{0},0,0,x_{3},0,0,0)$ is the zero form in $\mathbb{C}[x_{0},x_{3}]$, hence we get the conditon
@@
By Remark 2.4 the line $L$ is of type II can be rephrased in terms of the derivatives $\partial_{i}F|_{L}$, and we get the additional condition
@@
To conclude the proof we check with the aid of Macaulay2 that the surface defined by det$(M)$ for polynomials $Q_{0},\ Q_{1},$ and $P$ with coefficients satisfying conditions (2), (3), and (4) has a singularity of type $A_{3}$ at the point $q$. See ancillary file. ∎
@@
The ramification at the general point of the ramification locus of ${\widehat{\psi}}$ is simple.
This answers the natural question raised in [GK23a, Remark 3.13].
Proof.
Let $b$ be a general point in the branch divisor of $\psi$. Then by Theorem 2.5 the preimage ${\widehat{\psi}}^{-1}(b)$ consists of a finite number of points $p_{1},\ldots p_{k}$ where only $p_{1}$ is of ramification, while ${\widehat{\psi}}$ is étale at the points $p_{2},\ldots p_{k}$. All the points $p_{i}$ correspond to singularities of the quintic surface $S_{b}$: in particular $p_{2},\ldots p_{k}$ are singularities of type $A_{1}$, in contrast $p_{1}$ is a singular point of type $A_{3}$ by Theorem 2.6. The surface $S_{b}$ is specialization of the surface $S_{r}$ for the general line $r\in F$, which has exactly 16 nodes. Thus we deduce that 2 points among these 16 specialize to the singularity of type $A_{3}$, whereas the remaining 14 specialize to the $p_{2},..,p_{k}$, deducing that $S_{b}$ has 14 singularities of type $A_{1}$ apart from the $A_{3}$ singularity. ∎
$F_{r}:=\{\ell\in F:L\cap R\not=\emptyset\}.$
If $r$ is general, then $F_{r}$ is a smooth surface [Voi86, §3  Lemma 1]. As an application of our analysis we determine the singularities of $F_{r}$ for a general point $r$ in the branch locus of ${\widehat{\psi}}$.
Corollary 2.8.
Let $r\in F$ be a general line in the branch locus of ${\widehat{\psi}}$, then the surface $F_{r}$ has one singularity of type $A_{1}$ at the point $\ell$ corresponding to the unique line $L$ of type $\operatorname{II}$ with residual line $R$.
@@
$f\colon F_{r}\dashrightarrow S_{r},\quad\ell\mapsto\pi_{R}(L).$
The map clearly extends on the point $r$ by mapping $r$ to $\pi_{r}(\Lambda_{r})$, when the line $r$ is of type $I$. The map $f$ is 2-to-1 and can be regarded as the quotient of an involution, which, in case $r$ is general, it has 16 fixed points mapped to the 16 singular points of $S_{r}$. If $r$ is as in the hypothesis, then $S_{r}$ is has 14 singularities of type $A_{1}$ and one singularity of type $A_{3}$ corresponding to $\ell$. All the nodes get resolved in the 2-to-1 cover $F_{r}$, while the $A_{3}$ singularity gives an $A_{1}$ singularity, hence the claim. ∎

### v2

# Lines, Twisted Cubics on Cubic Fourfolds, and the Monodromy of the Voisin Map

@@ Abstract.
For a general cubic fourfold $Y$ with associated Fano variety of lines $F$, we show that the monodromy group of the finite degree 16 rational Voisin self-map $\psi\colon F\dashrightarrow F$ is maximal. To achieve this, we investigate the intriguing interplay between $\psi$ and the fixed locus of the antisymplectic involution on the LLSvS variety $Z$, examined via the degree 6 Voisin map $F\times F\dashrightarrow Z$.
@@ 1. Introduction
Let $Y$ be a smooth cubic fourfold and $F$ be its (Fano) variety of lines. The rich geometry of these varieties has attracted great attention from the mathematical community for several reasons. Notably, the Fano variety $F$ is one of the earliest examples of a locally complete family of projective hyperkähler manifolds. It is deformation equivalent to the Hilbert square of a K3 surface, and has natural polarisation of degree 6 and divisibility 2 [BD85]. The duo YFY-F between a Fano variety and a hyperkähler manifold has inspired numerous other constructions.
One distinguishing feature of $F$ is the degree 16 rational Voisin self-map $\psi\colon F\dashrightarrow F$. Remarkably, $\psi$ is the only known finite, non-birational self-map defined on a locally complete family of projective hyperkähler manifolds. Finite maps such as $\psi$ are particularly subtle to study as, unlike birational maps, they cannot be detected through their action on the second cohomology group and lack a comprehensive classification result as Hodge Torelli theorem that would allow precise control over their behavior. Since its construction [Voi04], this map has continued to attract significant interest in the mathematical community, as demonstrated by the numerous works on the subject[Ame09, GK23, GK25, GK22].
In this paper, we focus on the monodromy of the Voisin map $\psi$. The monodromy group is a discrete invariant that encodes the symmetries and intrinsic geometry of finite maps. Determining the monodromy group of branched coverings is a classical problem, originating with Jordan in the 1870s. This topic was revitalised by Harris, who provided a modern framework by proving that the monodromy and Galois groups coincide, and developed tools to establish when the monodromy group is maximal. Significant progress was made by Vakil, who introduced innovative techniques to study monodromy in Schubert problems on Grassmannians [Vak06].
Building on these developments, we take a step further by investigating linear spaces on a cubic hypersurface of dimension 4. Our main result shows that given a general cubic fourfold the monodromy group of the Voisin map $\psi$ is maximal, meaning that it is the full symmetric group on 16 elements. To achieve this, we leverage another Voisin map involving twisted cubics on the cubic fourfold, which unveils a fascinating connection with the fixed locus of the natural anti-symplectic involution acting on the LLSvS variety.
@@
$\Lambda_{L}:=\bigcap_{p\in L}T_{p}Y$
is 2-dimensional and in this case the line is said to be of type I. For special lines, called of type II, the dimension of $\Lambda_{L}$ is 3, see Proposition 2.1 and Definition 2.2 for more details. For the general line we have $\Lambda_{L}\cap Y=2L+R$ for a line $R\in F$, and one sets $\psi(L):=R$. This map has been studied in detail in [Ame09], its indeterminacy locus consists of the lines of type $\operatorname{II}$, which for a general cubic fourfold form a smooth surface $S_{\operatorname{II}}$. Blowing up $S_{\operatorname{II}}$ provides a resolution of the indeterminacy
@@
Gounelas and Kouvidakis recently computed that the restriction of $^{\widehat{\psi}}$ to the exceptional divisor $E$ of the blow-up morphism, which coincides with the ramification divisor of $^{\widehat{\psi}}$, is birational onto the image.
In Section 2 we study the map $^{\widehat{\psi}}$ by considering the projection of $Y$ from a general line $R$ in $Y$. Its resolution is a conic bundle over $\mathbb{P}^{3}$ with discriminant locus a quintic surface $S_{R}$ with exactly 16 nodes. These nodes correspond to the preimage of $R$ under $^{\widehat{\psi}}$. Nodal quintic surfaces have been classically studied by Beauville and Catanese [Bea80, Cat81] and more recently studied in [Huy24, BFK+24, CicwYCC+24]. First we study the singularities of $S_{R}$ for a special line $R$.
@@
Let $R\in F$ be a general line in the branch divisor of $^{\widehat{\psi}}$. Then the quintic surface $S_{R}$ has one singularity of type $A_{3}$ corresponding to the unique line $L$ of type $\operatorname{II}$ over $R$.
In Section 3 we shift our focus to the geometry of twisted cubics. With any smooth cubic fourfold $Y$ not containing a plane Lehn, Lehn, Sorger, and van Straten associated an eight-dimensional hyperkähler variety $Z$ parametrising families of twisted cubics and their flat degenerations. The variety $Z$ is equipped with a natural antisymplectic involution $\tau$ [Leh15]. Its fixed locus is a smooth Lagrangian submanifold with 2 connected components: one is isomorphic to the cubic fourfold $Y$, the other one, $W$, is of general type [FMOS23] and remains somewhat mysterious. Using the degree 6 rational map $\varphi\colon F\times F\dashrightarrow Z$ constructed by Voisin [Voi16], we offer an alternative description of $W$. We define the variety $P$ as the closure in $F\times F$ of
$\{(L_{1},L_{2})\in F\times F:L_{i}\text{ are of type I,\quad$L_{1}\not=L_{2}$,\quad and }\quad\psi(L_{1})=\psi(L_{2})\},$
which is birational to an irreducible component of the self-product of $^{\widehat{F}}$ over $F$.
@@
In Section 4 after revising the basic notions of monodromy we tackle the study of $^{\widehat{\psi}}$. Despite extensive study and numerous works concerning $^{\widehat{\psi}}$, for example about its entropy [Ame09], many of its properties remain elusive. Through an investigation of the restriction of the map $\varphi$ to the variety $P$ we prove the following
@@ Acknowledgments
This project began a long time ago, and over the years we benefited from conversations with many people. It is our pleasure to thank everybody who expressed interest and shared their point of view, especially Enrico Fatighenti, Frank Gounelas, Christian Lehn, Emanuele Macrì, Giovanni Mongardi, Alan Thompson, and Yilong Zhang. We also express our gratitude to the anonymous referees for their careful reading. Franco Giovenzana was funded by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Projektnummer 509501007, and partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC-2020- SyG-854361- HyperK). All authors are members of INdAM GNSAGA.
2. Nodal quintic surfaces with 16 nodes
In this section we recall basic facts about lines on cubic fourfolds and various properties of the Voisin map on $F$. Then we move on studying degenerations of nodal quintic surfaces with an even set of 16 nodes naturally arising in this context.
@@ 2. Nodal quintic surfaces with 16 nodes
Given a line $L$, we say that $L$ is a line of type I if the equivalent conditions (1)(4)(1)-(4) hold, whereas if (5)(8)(5)-(8) hold we say that $L$ is a line of type II. We set $\Lambda_{L}:=\cap_{P\in L}T_{P}Y$.
@@
In case $L\in F$ is of type $\operatorname{II}$, then the line $\mathscr{G}(L)\subset{\mathbb{P}}(V^{\vee})$ is the projective dual of $\Lambda_{L}\subset{\mathbb{P}}(V)$. In case of a line of type I, $\mathscr{G}(L)$ spans a ${\mathbb{P}}^{2}$ in ${\mathbb{P}}(V^{\vee})$ which is dual to $\Lambda_{L}$.
@@
[Huy23, Remark 2.2.2] Let $Y=V(f)\subset{\mathbb{P}}^{5}$ be a smooth cubic fourfold, let $L$ be a line in $Y$, then $L$ is of type II if and only if the partial derivatives $\partial_{0}f|_{L},...,\partial_{5}f|_{L}$ span a vector space of dimension 2 in $H^{0}(L,\mathcal{O}_{\mathbb{P}^{5}}(2))$.
Voisin observed that if $L$ is a line of type I, then $\Lambda_{L}\cap Y$ is a cubic curve, which, as it contains $L$ with multiplicity 2, consists of $L$ and a residual line $R$. As the general line is of type I, one gets the following
Definition 2.5.
The Voisin map $\psi$ is the rational map defined by mapping a line $L$ of type I to its residual $R$
$\psi\colon F\dashrightarrow F,\quad L\mapsto R.$
In general we say that the line $R$ is residual to $L$ if there exists a projective plane $\Lambda$ such that $\Lambda\cap Y=2L+R$. Following Gounelas and Kouvidakis if a line is residual to itself, we refer to it as triple line. For the general cubic fourfold the locus of triple lines is an irreducible surface of general type [GK22, Theorem A]. We denote it $V$.
The map $\psi$ has been studied in [Voi04] and [Ame09], where it is proven to be finite of degree 16. For a general cubic fourfold, lines of type II form a smooth surface $S_{\operatorname{II}}$ in $F$, and blowing up $F$ in this surface resolves the indeterminacy of $\psi$:
@@
The blow up $^{\widehat{F}}$ can be identified as the closure of the graph [GK22, Lemma 4.1],[Huy23, Remark 2.2.19].
$F\dashrightarrow\operatorname{Gr}(3,6),\ L\mapsto\Lambda_{L}.$
Elements in the exceptional locus are then just pairs $(L,\Xi)$, where $L$ is a line of type II and $\Xi$ is a projective plane such that $L\subset\Xi\subset\Lambda_{L}$. For such 2-dimensional spaces $\Xi$ we have that $\Xi\cap Y=2L+R$, in particular $R$ is residual to $L$.
@@
${{\widetilde{Y}}:=\operatorname{Bl}_{R}Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$$\scriptstyle{{\widetilde{\pi}}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{R}}$${{\mathbb{P}}^{3}}$
where $\pi_{R}$ is the projection from $R$ and $\widetilde{Y}$} is the blow-up of $Y$ in $R$. The morphism $\widetilde{\pi}$} is a conic bundle with discriminant a quintic surface $S_{R}$, whose singular locus consists of 16 nodes. The 16 nodes correspond to the preimage of $R$ under $^{\widehat{\psi}}$. Indeed, if $p_{i}$ denote the nodes of $S_{R}$ for $i=1,..,16$, then $L_{i}:=p({\widetilde{\pi}}^{-1}(p_{i}))$ are the 16 lines for which $\psi(L_{i})=R$ (see [Huy23, § 6.4.5] for an account on the various results in the literature about this surface).
Let $E$ be the exceptional divisor of the blow-up morphism ${\widehat{F}}\to F$. As $F$ has trivial canonical bundle, the divisor $E$ coincides with the ramification locus of the map $^{\widehat{\psi}}$.
Theorem 2.6 ([GK25, Theorem B] ).
Let $Y$ be a general cubic fourfold. The restriction ${\widehat{\psi}}|_{E}\colon E\to F$ is generically 1-to-1 onto the image.
Proposition 2.7.
Let $Y$ be a general cubic fourfold. The ramification at the general point of the ramification locus of $^{\widehat{\psi}}$ is simple.
Proof.
Comparing the pullbacks of the canonical of $F$ in diagram (5) one gets:
$E=pr^{*}(K_{F})+E=K_{\widehat{F}}=\widehat{\psi}^{*}(K_{F})+R=R$
where $R$ is the ramification locus with the natural schematic structure. Hence the ramification is simple by [BHPVdV04, Lemma I.16.1]
In other words, over the general point $R$ in the branch divisor of $^{\widehat{\psi}}$, there is exactly one point of ramification, i.e. there exists exactly one line $L$ of type II with residual $R$, meaning that ${\widehat{\psi}}(L,\langle L,R\rangle)=R$.
@@
Let $R\in F$ be a general line in the branch divisor of $^{\widehat{\psi}}$. Then the quintic surface $S_{R}$ has exactly 14 singularities of type $A_{1}$ and exactly 1 singularity of type $A_{3}$ corresponding to the unique line $L$ of type $\operatorname{II}$ with residual line $R$.
@@
Let $R$ be a general line in the branch locus of $^{\widehat{\psi}}$ and let $L$ be the unique line of type $\operatorname{II}$ with residual $R$. We may assume that $R$ is of type I and that is given as the zero set $V(x_{2},x_{3},x_{4},x_{5})$. Following Clemens and Griffiths [CG72, equation 6.9] we write the equation of the cubic fourfold as
(6) $f:=x_{4}x_{0}^{2}+x_{5}x_{0}x_{1}+x_{3}x_{1}^{2}+x_{0}Q_{0}+x_{1}Q_{1}+P=0,$
@@
We assume that the line $L$ is given by $V(x_{1},x_{2},x_{4},x_{5})$, then $p:=(1:0:0:0:0:0)$ is the intersection of $L$ and $R$, and $q:=(0:0:0:1:0:0)$ is the point of intersection with $\mathbb{P}^{3}$.
@@
The cubic fourfold $Y$ contains the line $L$ if $f(x_{0},0,0,x_{3},0,0,0)$ is the zero form in $\mathbb{C}[x_{0},x_{3}]$, hence we get the condition
@@
By Remark 2.4 the fact that the line $L$ is of type II can be rephrased in terms of the derivatives $\partial_{i}f|_{L}$, and we get the additional condition
@@
To conclude the proof we check with the aid of Macaulay2 that the surface defined by det(M)(M) for polynomials $Q_{0},\ Q_{1},$ and $P$ with coefficients satisfying conditions (7), (8), and (9) has a singularity of type $A_{3}$ at the point $q$. See ancillary file. ∎
As a consequence we obtain an alternative proof of the simpleness of the ramification of $\widehat{\psi}$.
Proof.
(of Proposition 2.7) Let $b$ be a general point in the branch divisor of $\psi$. Then by Theorem 2.6 the preimage ${\widehat{\psi}}^{-1}(b)$ consists of a finite number of points $p_{1},\ldots p_{k}$ where only $p_{1}$ is of ramification, while $^{\widehat{\psi}}$ is étale at the points $p_{2},\ldots p_{k}$. All the points $p_{i}$ correspond to singularities of the quintic surface $S_{b}$: in particular $p_{2},\ldots p_{k}$ are singularities of type $A_{1}$, in contrast $p_{1}$ is a singular point of type $A_{3}$ by Theorem 2.8. The surface $S_{b}$ is a specialization of the surface $S_{R}$ for the general line $R\in F$, which has exactly 16 nodes. Thus we deduce that 2 points among these 16 specialize to the singularity of type $A_{3}$, whereas the remaining 14 specialize to the $p_{2},..,p_{k}$, deducing that $S_{b}$ has 14 singularities of type $A_{1}$ apart from the $A_{3}$ singularity. ∎
$F_{R}:=\{L\in F:L\cap R\not$$\emptyset\}$.
If $R$ is general, then $F_{R}$ is a smooth surface [Voi86, §3  Lemma 1]. As an application of our analysis we study the singularities of $F_{R}$ for a general point $R$ in the branch locus of $^{\widehat{\psi}}$.
@@
Let $R\in F$ be a general line in the branch locus of $^{\widehat{\psi}}$. If the surface $F_{R}$ is normal, then it has exactly one singular point, which is of type $A_{1}$ and corresponds to the unique line $L$ of type $\operatorname{II}$ with residual line $R$.
@@
$f\colon F_{R}\dashrightarrow S_{R},\quad L\mapsto\pi_{R}(L).$
The map clearly extends to the point $R$ by mapping $R$ to $\pi_{R}(\Lambda_{R})$, when the line $R$ is of type II. The map $f$ is 2-to-1 and can be regarded as the quotient of an involution, which, in case $R$ is general, has 16 fixed points mapped to the 16 singular points of $S_{R}$. If $R$ is as in the hypothesis, then the fixed points of this involution are in bijection with the singularities of $S_{R}$: 14 singularities of type $A_{1}$ and one singularity of type $A_{3}$ corresponding to the unique line $L$ of type $\operatorname{II}$ with residual $R$ by Theorem 2.8.
As $F_{R}$ is normal, the morphism $f$ around each singular point is the unique morphism characterized by [GKP16, Proposition 3.13]. Locally around the $A_{1}$ singular point it is isomorphic to the quotient presentation $\mathbb{C}^{2}\to\mathbb{C}^{2}/\mu_{2}$ of the $A_{1}$ singularity, and around the $A_{3}$ singular point to the presentation $\mathbb{C}^{2}/\mu_{2}\to\mathbb{C}^{2}/\mu_{4}$ of the $A_{3}$ singularity as quotient of the $A_{1}$ singularity. ∎